# Temporally Compressed 3D Gaussian Splatting for Dynamic Scenes

### v1

@@ Abstract
footnotetext: Equal Contribution.
Recent advancements in high-fidelity dynamic scene reconstruction have leveraged dynamic 3D Gaussians and 4D Gaussian Splatting for realistic scene representation. However, to make these methods viable for real-time applications such as AR/VR, gaming, and rendering on low-power devices, substantial reductions in memory usage and improvements in rendering efficiency are required. While many state-of-the-art methods prioritize lightweight implementations, they struggle in handling scenes with complex motions or long sequences. In this work, we introduce Temporally Compressed 3D Gaussian Splatting (TC3DGS), a novel technique designed specifically to effectively compress dynamic 3D Gaussian representations. TC3DGS selectively prunes Gaussians based on their temporal relevance and employs gradient-aware mixed-precision quantization to dynamically compress Gaussian parameters. It additionally relies on a variation of the Ramer-Douglas-Peucker algorithm in a post-processing step to further reduce storage by interpolating Gaussian trajectories across frames. Our experiments across multiple datasets demonstrate that TC3DGS achieves up to 67$\times$ compression with minimal or no degradation in visual quality.
Figure 1: Our method, TC3DGS, achieves up to 67× compression with minimal loss in image quality, while demonstrating significantly higher rendering speed compared to [26].
1 Introduction
Dynamic scene reconstruction is essential for applications in virtual and augmented reality, gaming and robotics, where a real-time and accurate representation of moving objects and their environment is key to immersive experiences. Recent advancements such as Neural Radiance Fields (NeRF) [28] have enabled high-fidelity scene generation, albeit at the cost of slow rendering speeds. To address this limitation, 3D Gaussian Splatting (3DGS) [20] leverages sparse Gaussian splats for efficient scene rendering, particularly for static scenes.
Since the advent of 3DGS, a plethora of extensions to dynamic scenes have been proposed [39, 26, 6, 40, 41, 8]. In this context, some methods [39, 26, 6, 40] allow the Gaussians to evolve over time, capturing time-varying properties such as position, opacity, and covariance. In contrast, other methods learn spatio-temporal Gaussians to represent the dynamic content directly, which allows for a more flexible modeling of scene variations [6, 40, 8, 41]. However, obtaining a high-quality representation with these methods often requires a large number of Gaussians, leading to substantial storage overhead. Furthermore, as shown in Figure 2, spatio-temporal methods face the challenge of effectively adjusting the Gaussian parameters when applied to dynamic scenes with rapid and complex motions, such as those in [18]. This is alleviated by Dynamic 3D Gaussians [26], which enforces consistency across all frames, but at the cost of increased storage and rendering time.
To address these challenges, we propose Temporally Compressed 3D Gaussian Splatting (TC3DGS), a novel approach designed to efficiently compress dynamic 3D Gaussian representations for high-quality, real-time scene rendering. Unlike traditional methods, TC3DGS reduces both the number and the memory footprint of the Gaussians, by selectively pruning splats based on temporal importance and learning a parameter-specific bit-precision. This selective compression allows us to maintain scene fidelity while significantly reducing storage and computational requirements, making TC3DGS well-suited for dynamic environments with complex motions.
Our approach begins with a pruning and masking strategy designed to eliminate redundant Gaussians. While previous works [42, 11, 21] have proposed pruning methods for 3DGS, they do not attempt to model dynamic scenes and thus take no advantage of temporal compression. Unfortunately, adapting these pruning strategies to dynamic scenes is not straightforward, as different Gaussians may need to be pruned across different frames. Here, we introduce a method that explicitly integrates this temporal aspect into the training objective, allowing us to prune dynamic Gaussian splats more effectively. To further optimize memory usage in our scene representation, we also aim to compress the storage size of the remaining Gaussians. To this end, we developed a gradient-aware mixed-precision quantization method that adjusts the bit precision of each Gaussian parameter based on its sensitivity. We use gradient magnitudes to determine parameters with high sensitivity, and allocate them a higher precision, while those with lower impact on the scene are quantized with fewer bits. Therefore, our method achieves a good balance between compression and reconstruction accuracy. Finally, to further enhance the efficiency of our dynamic representation, we introduce a keypoint extraction algorithm as a post-processing step. This algorithm simplifies the temporal trajectories of the Gaussian parameters, preserving only the key data points that are crucial for accurate dynamic scene representation. This greatly reduces the amount of temporal data stored, allowing for more compact scene representations, as demonstrated in our experiments where we achieve a compression rate of up to 67$\times$ while preserving rendering quality.
Our experiments on benchmark datasets demonstrate the benefits of our method across diverse scenarios. Furthermore, we perform an ablation study to demonstrate the contribution of each key component in our pipeline. In a nutshell, our major contributions are as follows:
  • We develop a sensitivity-driven, gradient-based quantization method that dynamically assigns bit precision to parameters based on their impact on reconstruction accuracy, optimizing memory usage.
@@ 1 Introduction
Ground Truth Dynamic 3DGS [26] TC3DGS (Ours) STG [23] 4D Gaussian [38]
Figure 2: Comparative Evaluation on Panoptic Dataset [18] Recent lightweight state-of-the-art methods (STG [23], and 4D-Gaussian [38]) struggle to accurately reconstruct scenes in complex environments. In contrast, our compression strategy effectively captures and maintains high fidelity in dynamic scene details.
@@ 2 Related Work
Recent advancements in 3D reconstruction based on Neural Radiance Fields (NeRFs) [28] and 3D Gaussian Splatting (3DGS) [20] have achieved remarkable levels of visual fidelity and accuracy. These methods have subsequently been extended to 4D representations [33, 24, 13], enabling dynamic scene reconstruction. Methods to decompose a 4D scene into multiple 2D planes to learn a more compact representation are also explored by various methods [16, 3, 35, 1]. In the case of 3DGS, where Gaussians are explicitly stored and rendered, different approaches have emerged to model their time dependence. One prominent line of work [39, 26, 6, 40] in this area optimizes a canonical set of Gaussians from the initial frame, and combines it with a deformation motion field allowing temporal variations of the Gaussian parameters. However, these methods are limited to short videos, as they cannot add Gaussians after the initial frame.
Another class of methods [41, 8, 19, 9] directly models temporal Gaussians that can be present in a subset of frames, enabling certain elements to appear in selected time ranges and thus increasing the expressivity of their reconstruction. However, a major limitation across these methods is that both training and inference times for novel view synthesis scale with the number of Gaussians, the length of the sequence, and the complexity of their parameters, presenting a key bottleneck in enhancing reconstruction quality.
@@
An important research direction has thus emerged in developing more compact representations of radiance fields. For NeRFs, compact grid structures [30, 4, 15, 14] have already proven effective in reducing network sizes and enhancing accuracy.
@@
With 3DGS, recent works have either concentrated on optimizing the representation of the Gaussian parameters [11], or on identifying low-importance Gaussians and pruning them entirely [42, 11, 21]. Unorthodox techniques like representing the Gaussian parameters as 2D grids and applying image compression techniques [29, 31] have also been studied. Better initialization and weighted sampling based pruning [12] has also shown promising results. Additionally, the use of traditional compression techniques such as vector quantization [21, 31] or entropy models [5] have shown some potential for the compression of static scenes, but scaling them to dynamic scenes with possibly hundreds or thousands of frames remains a challenge.
Indeed, dynamic scenes require an even larger set of parameters to accurately capture motion, temporal variations, and complex interactions within the scene.
@@
The temporal relevance of each Gaussian changes dynamically, and traditional pruning strategies designed for static scenes are insufficient, as they lack adaptability to these fluctuations. To the best of our knowledge, we are the first to propose a compression framework specifically designed for dynamic 3D Gaussians. Our approach combines temporal relevance-based pruning, gradient-based mixed-precision quantization, and trajectory simplification to address the unique requirements of dynamic scenes.
3 Method
@@
In this section, we will first briefly discuss the Dynamic 3DGS [26] method in Sec. 3.1, which serves as the foundation of our proposed approach. We then detail in Sec. 3.2.1 our novel masking and pruning strategies designed to eliminate redundant Gaussians. Next, we present a sensitivity-based mixed precision technique for efficient parameter compression in Sec. 3.2.2, followed by a post-training compression strategy aimed at minimizing storage overhead in Sec. 3.2.3.
@@
Figure 3: Overview of our Temporally Compressed 3D Gaussian Splatting for Dynamic Scenese (TC3DGS) method. Our approach involves a temporally consistent masking strategy to select relevant 3D Gaussians across frames. The masked Gaussians are then pruned and quantized using a gradient-based, parameter-aware bit-precision quantization scheme, which optimizes memory usage while preserving scene fidelity.
3.1 Dynamic 3D Gaussians
Dynamic 3DGS [26] models dynamic scenes by allowing the Gaussians to move and rotate over time while enforcing that they have persistent color, opacity, and scale. This approach reconstructs a dynamic 3D scene over time using a series of images taken from multiple cameras across different time steps, along with the cameras’ intrinsic and extrinsic parameters. The method sequentially reconstructs a dynamic 3D scene by initializing each time step from the previous one and performing test-time optimization without additional training data. This is unlike most other concurrent methods, which attempt to optimize all frames jointly. This conceptual difference allows Dynamic 3DGS [26] to enforce consistency in the reconstructed Gaussians across frames. Additionally, this method enables tracking objects throughout the dynamic scene, as each 3D Gaussian has a unique correspondence across frames.
Each dynamic scene $\mathcal{S}$ is parameterized by a set of Dynamic 3D Gaussians, with certain parameters fixed for the entire sequence and others varying over time. Specifically, each Gaussian retains a fixed 3D scale $(s_{x},s_{y},s_{z})$, color $(r,g,b)$, opacity logit $o$, and background logit $bg$ as determined in the initial frame. In contrast, parameters such as the 3D center $(x_{t},y_{t},z_{t})$ and 3D rotation, expressed as a quaternion $(q_{w,t},q_{x,t},q_{y,t},q_{z,t})$, evolve over time. This formulation enables the Gaussians to represent consistent scene elements across frames while dynamically adjusting their positions and orientations.
During training, this 3D scene representation is iteratively updated using a differentiable renderer to minimize the photometric error between the rendered and input images. Specifically, each Gaussian influences a point $p$ in physical 3D space according to the standard (unnormalized) Gaussian equation weighted by its opacity, i.e.,
$f_{i,t}(p)$ $=\sigma({o_{i}})\exp\left(-\frac{1}{2}(p-\mu_{i,t})^{T}\Sigma_{i,t}^{-1}(p-\mu_{i,t})\right)\;,$
[…]
@@ 3.2 TC3DGS
Dynamic 3DGS [26] is particularly promising because it models the dynamic scene as movements of Gaussians under kinematic constraints w.r.t. the previous timestep. By optimizing the position and rotation of the Gaussians instead of learning deformation functions, it removes the limitation on possible deformations due to the characteristics of the modeling function. However, by learning position and rotation at each timestep separately, the number of parameters increases linearly, resulting in large storage sizes, and thus limiting its applicability to high-fidelity and long-range scene modeling.
@@ 3.2.1 Gaussian Masking and Pruning
Pruning techniques aimed at identifying and removing low-importance Gaussians have been successfully applied to static 3D Gaussian splatting [21, 32, 17, 11]. However, for dynamic scenes, a temporally consistent importance measure is required to ensure that the pruned Gaussians remain insignificant throughout the scene’s duration.
Various approaches to computing the importance of each Gaussian in static scenes using training images and camera positions have bee proposed [32, 11]. These methods focus on the contribution of each Gaussian to the training views. In dynamic scenes, Gaussians are not stationary, so their contributions vary over time. To prune Gaussians effectively from dynamic scenes, it is essential to maintain the contributions of high-importance Gaussians consistently high, while suppressing low-importance ones, thus enabling more effective pruning.
Compact-3DGS [21] introduces a masking approach based on Gaussian volume and transparency. Gaussians with low opacity, minimal volume, or both are masked out since they have negligible impact on the rendered images. A mask parameter $m\in\mathbb{R}^{N}$ is learned to produce binary masks $M\in\{0,1\}^{N}$ using a straight-through estimator. This binary mask is then applied to the Gaussians by scaling their opacities and sizes. This is expressed as
@@
where $n$ represents the Gaussian index, $\epsilon$ denotes the masking threshold, $\operatorname*{sg}(\cdot)$ is the stop-gradient operator, and $\mathds{1}[\cdot]$ and $\sigma(\cdot)$ correspond to the indicator and sigmoid functions, respectively.
Compact-3DGS [21] explains that this mask learns to remove Gaussians with low opacity and/or small volume. However, as shown in Figure 4, this explanation does not hold in dynamic scenes. The value of $m_{n}$ can change significantly across time frames, even with fixed opacity $o$ and scale $s$ after the initial time frame.
@@
Since the values in $m$ strictly need to be high due to their role in rendering through Eq. 1 and the photometric loss, we introduce an additional regularization to incentivize lower values. We thus regularize $m$ by minimizing
@@
A key motivation for ensuring consistency of $m_{n,t}$ across frames is to capture the global importance of the Gaussians during pruning. To achieve this, we learn $m_{n,t}$ for each timestamp and introduce a consistency loss function that encourages $m_{n,t}$ to remain close to $m_{n,t-1}$. Specifically, we define this mask consistency loss function as

### v2

# Lausanne, Switzerland

@@ Abstract
[…]
Project Page: https://ahmad-jarrar.github.io/tc-3dgs/
@@
Ground Truth D-3DGS [luiten2023dynamic] TC3DGS (Ours) STG [stg] 4D Gaussian [4dgs]
Figure 1: Comparative Evaluation on Panoptic Dataset [Joo_2019] Recent lightweight state-of-the-art methods (STG [stg], and 4D-Gaussian [4dgs]) struggles to reconstruct scenes in complex environments accurately. In contrast, our compression strategy effectively captures and maintains high fidelity.
1 Introduction
Dynamic scene reconstruction is essential for applications in virtual and augmented reality, gaming, and robotics, where a real-time and accurate representation of moving objects and their environment is the key to immersive experiences. Recent progress such as Neural Radiance Fields (NeRF) [mildenhall2021nerf] has enabled high-fidelity scene generation, although at the cost of a slow rendering speed. To address this limitation, 3D Gaussian Splatting (3DGS) [kerbl20233d] leverages sparse Gaussian splats for efficient scene rendering, particularly for static scenes.
Since the advent of 3DGS, various methods have been proposed to extend dynamic scene modeling, either by evolving Gaussians over time to capture time-varying properties [4dgs, luiten2023dynamic, das2024neural, yang2024deformable, xu20234k4d] or by learning spatio-temporal Gaussians for more flexible scene representation [das2024neural, yang2024deformable, duan20244d, yang2023real, Jiang_2024_CVPR]. Our experiments with such frameworks, illustrated in Figure 1 and in supplementary videos have highlighted that many state-of-the-art methods struggle to effectively adjust Gaussian parameters in dynamic scenes with rapid and complex motions, such as those of [Joo_2019]. The only exception is Dynamic 3D Gaussians [luiten2023dynamic], which mitigates this issue by enforcing consistency across frames. However, this comes with a dramatic increase in storage and rendering costs, making this method impractical for real-world applications, such as deployment on VR/AR headsets.
To address these challenges, we propose Temporally Compressed 3D Gaussian Splatting (TC3DGS), a novel approach designed to efficiently compress dynamic 3D Gaussian representations for high-quality, real-time scene rendering. Unlike traditional methods, TC3DGS reduces both the number and the memory footprint of the Gaussians by selectively pruning splats based on temporal importance and learning a parameter-specific bit-precision. This selective compression allows us to maintain scene fidelity while significantly reducing storage and computational requirements, making TC3DGS well-suited for dynamic environments with complex motions.
[…]
Our experiments on benchmark datasets demonstrate the benefits of our method on diverse scenarios. Furthermore, we perform an ablation study to demonstrate the contribution of each key component in our pipeline. In a nutshell, our major contributions are as follows:
  • We identify and analyze the failure of existing methods in handling scenes with rapid and complex motion, highlighting these shortcomings through qualitative and quantitative evaluations.
  • We develop a sensitivity-driven, gradient-based mixed-precision quantization method that dynamically assigns bit precision to parameters based on their impact on reconstruction accuracy, optimizing memory usage.
@@ 2 Related Work
Recent advancements in 3D reconstruction based on Neural Radiance Fields (NeRFs) [mildenhall2021nerf] and 3D Gaussian Splatting (3DGS) [kerbl20233d] have achieved remarkable levels of visual fidelity and accuracy. These methods have subsequently been extended to 4D representations [pumarola2021d, liu2023robust, TiNeuVox, kulhanek2024wildgaussians, Li2024ST], enabling dynamic scene reconstruction. Methods to decompose a 4D scene into multiple 2D planes to learn a more compact representation are also explored by various methods [kplanes_2023, Cao2023HexPlane, shao2023tensor4d, attal2023hyperreel]. In the case of 3DGS, where Gaussians are explicitly stored and rendered, different approaches have emerged to model their time dependence. One prominent line of work [4dgs, luiten2023dynamic, das2024neural, yang2024deformable, huang2024scgs, kratimenos2024dynmf] in this area optimizes a canonical set of Gaussians from the initial frame, and combines it with a deformation motion field allowing temporal variations of the Gaussian parameters. However, these methods are limited to short videos, as they cannot add Gaussians after the initial frame.Recent works [sun20243dgstream, hicom2024] have explored online causal training of 3DGS models for streaming applications and to handle long scenes. However, these methods scale with the number of frames, like Dynamic 3DGS [luiten2023dynamic].
Another class of methods [yang2023real, duan20244d, katsumata2025compact] directly models temporal Gaussians that can be present in a subset of frames, enabling certain elements to appear in selected time ranges and thus increasing the expressivity of their reconstruction. However, a major limitation across these methods is that both training and inference times for novel view synthesis scale with the number of Gaussians, the length of the sequence, and the complexity of their parameters, presenting a key bottleneck in enhancing reconstruction quality.
@@
An important research direction has thus emerged in developing more compact representations of radiance fields. For NeRFs, compact grid structures [muller2022instant, chen2022tensorf, fridovich2022plenoxels, franke2024trips, scaffoldgs] have already proven effective in reducing network sizes and enhancing accuracy.
@@
With 3DGS, recent works have either concentrated on optimizing the representation of the Gaussian parameters [fan2023lightgaussian, chen2024fast], or on identifying low-importance Gaussians and pruning them entirely [zhang2024lp, fan2023lightgaussian, lee2024compact, girish2024eaglesefficientaccelerated3d]. Unorthodox techniques like representing the Gaussian parameters as 2D grids and applying image compression techniques [morgenstern2024self_grid, navaneet2024compgs] have also been studied. Better initialization and weighted sampling based pruning [Fang2024MiniSplattingRS] has also shown promising results. Additionally, the use of traditional compression techniques such as vector quantization [lee2024compact, navaneet2024compgs] or entropy models [chen2025hac] have shown some potential for the compression of static scenes, but scaling them to dynamic scenes with possibly hundreds or thousands of frames remains a challenge.
@@
Indeed, dynamic scenes require an even larger set of parameters to accurately capture motion, temporal variations, and complex interactions within the scene. The temporal relevance of each Gaussian changes dynamically, and traditional pruning strategies designed for static scenes are insufficient, as they lack adaptability to these fluctuations. To the best of our knowledge, we are the first to propose a compression framework specifically designed for dynamic 3D Gaussians. Our approach combines temporal relevance-based pruning, gradient-based mixed-precision quantization, and trajectory simplification to address the unique requirements of dynamic scenes.
3 On the Challenges of Fast Motion
One of the most widely used datasets for benchmarking dynamic novel view synthesis, the Neural 3D Video dataset [li2022neural3dvideosynthesis], consists exclusively of indoor scenes with objects moving within their local vicinity. Existing methods thus perform well, as object motion remains limited. However, when we applied these methods to the Panoptic Sports dataset [Joo_2019], their performance deteriorated significantly due to the presence of fast moving objects that travel longer distances. Below, we analyze the underlying reasons why these methods struggle on such complex datasets.
Existing Gaussian Splatting methods for dynamic scenes often incorporate motion priors to regulate the movement of the Gaussian points. These priors determine the type of motion for which the method is more suitable. The key distinction between different approaches is how they integrate these motion priors.
Specifically, deformation field-based methods, such as 4D-Gaussians [4dgs] and E-D3DGS [bae2024ed3dgs], implicitly encode motion priors within the structure of neural networks, yielding flexible but less interpretable motion representations. These methods fails in the presence of large sudden movements due to the spectral bias of neural networks towards simpler functions [rahaman2019spectral].
As an alternative, STG [stg] models point trajectories as polynomial curves. However, this approach restricts the representation to 3rd degree polynomials, which limits the types of motion it can represent. While higher-degree polynomials enable the representation of more complex motion, they also introduce significant storage overhead.
Deformable 3DGS [yang2024deformable] impose explicit constraints, e.g., local rigidity, to regulate motion in a more interpretable manner. However, the use of a neural network in Deformable 3DGS to model the deformation field inherits the limitations of 4D-Gaussians and E-D3DGS discussed earlier.
Among the methods discussed above only Dynamic 3DGS [luiten2023dynamic], deformable field and explicit constraints based model, is trained causally, i.e., learning Gaussian deformations frame by frame, instead of optimizing the Gaussians across all time frames simultaneously, as done by the other methods. This causal training strategy requires storing and optimizing point positions in every frame, offering greater flexibility in motion modeling despite the local rigidity constraint, as the Gaussians are optimized independently for each frame. However, this comes at a high storage and rendering cost. Therefore, in section˜4, we propose a method to significantly compress Dynamic 3DGS while preserving its ability to capture complex motions.
In Figure 1, we provide a visual comparison illustrating how existing methods struggle with scenes featuring rapid and complex motion. In addition, we provide videos in the supplementary material for a more extensive evaluation.
4 Compression
@@
In this section, we introduce our TC3DGS approach, encompassing novel masking and pruning strategies designed to eliminate redundant Gaussians, a sensitivity-based mixed precision technique for efficient parameter compression, and a post-training compression strategy to minimize storage overhead. The overview of our method can be visualized in Figure 2.
@@
Figure 2: Overview of our Temporally Compressed 3D Gaussian Splatting for Dynamic Scenese (TC3DGS) method. Our approach involves a temporally consistent masking strategy to select relevant 3D Gaussians across frames. The masked Gaussians are then pruned and quantized using a gradient-based, parameter-aware bit-precision quantization scheme.
@@ 4.1 TC3DGS
Dynamic 3DGS [luiten2023dynamic] is particularly promising because it models the dynamic scene as movements of Gaussians under kinematic constraints w.r.t. the previous timestep. By optimizing the position and rotation of the Gaussians instead of learning deformation functions, it removes the limitation on possible deformations due to the characteristics of the modeling function. However, by learning position and rotation at each timestep separately, the number of parameters increases linearly, resulting in large storage sizes, and thus limiting its applicability to high-fidelity and long-range scene modeling.
@@ 4.1.1 Gaussian Masking and Pruning
Pruning techniques aimed at identifying and removing low-importance Gaussians have been successfully applied to static 3D Gaussian splatting [lee2024compact, Niedermayr_2024_CVPR, girish2024eaglesefficientaccelerated3d, fan2023lightgaussian]. However, for dynamic scenes, a temporally consistent importance measure is required to ensure that the pruned Gaussians remain insignificant throughout the scene duration.
Various approaches to computing the importance of each Gaussian in static scenes using training images and camera positions have bee proposed [Niedermayr_2024_CVPR, fan2023lightgaussian]. These methods focus on the contribution of each Gaussian to the training views. In dynamic scenes, Gaussians are not stationary, so their contributions vary over time. To prune Gaussians effectively from dynamic scenes, it is essential to maintain the contributions of high-importance Gaussians consistently high, while suppressing low-importance ones, thus enabling more effective pruning.
Compact-3DGS [lee2024compact] introduces a masking approach based on Gaussian volume and transparency. Gaussians with low opacity, minimal volume, or both are masked out because they have negligible impact on the rendered images. A mask parameter $m\in\mathbb{R}^{N}$ is learned to produce binary masks $M\in\{0,1\}^{N}$ using a straight-through estimator. This binary mask is then applied to the Gaussians by scaling their opacities and sizes. This is expressed as
@@
where $n$ represents the Gaussian index, $\epsilon$ denotes the masking threshold, $\operatorname*{sg}(\cdot)$ is the stop-gradient operator, $s_{n}\in\mathbb{R}^{3}$ represents scale of the gaussian, $o_{n}\in\mathbb{R}$ represents opacity of the gaussian, and $\mathds{1}[\cdot]$ and $\sigma(\cdot)$ correspond to the indicator and sigmoid functions.
In [lee2024compact], it is explained that this mask learns to remove Gaussians with low opacity and/or small volume. However, as shown in Figure 3, this explanation does not hold in dynamic scenes. The value of $m_{n}$ can change significantly across time frames, even with fixed opacity $o$ and scale $s$ after the initial time frame.
@@
Since the values in $m$ strictly need to be high due to their role in rendering through eq.˜1 and the photometric loss, we introduce an additional regularization to incentivize lower values. We thus regularize $m$ by minimizing
@@
A key motivation for ensuring consistency of $m_{n}$ across frames is to capture the global importance of the Gaussians during pruning. To achieve this, we replace $m_{n}$ with a time dependent $m_{n,t}$ defined for every timestamp separately and use it instead of $m_{n}$ in the rendering. We introduce a consistency loss function that encourages $m_{n,t}$ to remain close to $m_{n,t-1}$. Specifically, we define this mask consistency loss function as